OpenStudy (anonymous):

(sqrtx^2+24) - (sqrt13x-12) = 0

OpenStudy (anonymous):

Do you mean \[\sqrt{x^{2}+24}-\sqrt{13x-12}=0?\]

OpenStudy (anonymous):

If so, add \[\sqrt{13x-12}\] to both sides, square both sides, subtract 13x, add 12, factor(use quadratic if you can't factor) and solve.